### v3

# SCB-Dataset: A Dataset for Detecting Student Classroom Behavior

Abstract
Using deep learning methods to detect the classroom behaviors of both students and teachers is an effective way to automatically analyze classroom performance and enhance teaching effectiveness. Then, there is still a scarcity of publicly available high-quality datasets on student-teacher behaviors. Based on the SCB-Dataset3 we proposed previously, we have introduced a larger, more comprehensive, and higher-quality dataset of student-teacher classroom behaviors, known as SCB-Dataset5. Our dataset comprises 7428 images and 106830 labels across 20 classes: hand-raising, read, write, bow head, turn head, talk, guide, board writing, stand, answer, stage interaction, discuss, clap, yawn, screen, blackboard, teacher, leaning on the desk, using the phone, using the computer. We evaluated the dataset using the YOLOv7 series of algorithms We believe that SCB-Dataset5 can provide a solid foundation for future applications of artificial intelligence in education. Our SCB-Dataset5 can be downloaded at the following link: :https://github.com/Whiffe/SCB-dataset
Index Terms:
Keywords: deep learning, YOLO, student-teacher classroom behavior, SCB-Dataset5
I Introduction
In this study, we have iteratively optimized our previous work further to expand the SCB-Dataset[1]. Initially, we focused solely on the behavior of students raising their hands, but now we have expanded to include 20 classes: hand-raising, read, write, bow head, turn head, talk, guide, board writing, stand, answer, stage interaction, discuss, clap, yawn, screen, blackboard, teacher, leaning on the desk, using the phone, using the computer. This work further addressed the research gap in detecting student behaviors in classroom teaching scenarios.
Figure 1: An example from the SCB-dataset.
Existing student classroom behavior detection algorithms can be roughly divided into three categories: video-action-recognition-based[2], pose-estimation-based[3] and object-detection-based[4]. Video-based student classroom behavior detection enables the recognition of continuous behavior, which requires labeling a large number of samples. For example, the AVA dataset[5] for SlowFast[6] detection is annotated with 1.58M. And, video behavior recognition detection is not yet mature, as in UCF101[7] and Kinetics400[8], some actions can sometimes be determined by the context or scene alone. Pose-estimation-based algorithms characterize human behavior by obtaining position and motion information of each joint in the body, but they are not applicable for behavior detection in overcrowded classrooms. Considering the challenges at hand, object-detection-based algorithms present a promising solution. In fact, in recent years object-detection-based algorithms have made tremendous breakthroughs, such as YOLOv7[9]. Therefore, we have employed an object-detection-based algorithm in this paper to analyze student behavior.
As for object detection, the two-stage and one-stage object detection frameworks[10][11] have received more attention due to their impressive detection results on public datasets. However, the datasets from real classrooms are quite different from public ones and the classical methods perform poorly in real classrooms. One of the representative issues is large scale variations among different positions, such as students in the front row of the classroom (about 40×40 pixels) and students in the back row (about 200×200 pixels), which results in high scale variations of almost 25 times. To make matters worse, compared to the most popular object detection dataset MS COCO[12], The occlusion between students is very serious. Moreover, Behaviors vary greatly in different environments, among different people, and from different angles.
In this work, we explore the effectiveness of computer vision techniques in automatically analyzing student behavior patterns in the classroom. The dataset fills a gap in current research on detecting student behavior in classroom teaching scenarios.
We have conducted extensive data statistics and benchmark tests to ensure the dataset’s quality, providing reliable training data.
Our main contributions are as follows:
1. We have updated the SCB-Dataset to its fifth version (SCB-Dataset5), increasing to 20classes. This dataset contains a total of 7428 images with 106830 labels. It covers different scenarios from kindergarten to university.
2. We conducted extensive benchmark testing on the SCB-Dataset5, providing a solid foundation for future research.
3. We conducted in-depth multi-dimensional analysis and testing on the SCB-Dataset5, and successfully applied it to real classroom environments. Based on feedback from the school, we continue to iterate and optimize the SCB-Dataset5 to meet practical needs.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: Examples of images from various classes in the SCB dataset.
II Related Works
II-A Student classroom behavior dataset
In recent years, many researchers have adopted computer vision technology to automatically detect students’ classroom behaviors, but the lack of open student behavior data set in the field of education has severely limited the application of video behavior detection in this field. Many researchers have also proposed many unpublished datasets.
ActRec-Classroom[13] comprises 5 classes: listening, fatigue, raising hand, leaning, and reading/writing with 5,126 images. The method involves first using Faster R-CNN to detect human bodies, followed by OpenPose to extract key points of skeletons, faces, and fingers. Lastly, a CNN-based classifier is developed for action recognition.
A large-scale dataset for student behavior [15] , compiled from thirty schools, labels student behaviors using bounding boxes frame-by-frame and contains 70,000 hand-raising samples, 20,000 standing samples, and 3,000 sleeping samples. An enhanced Faster R-CNN model for student behavior analysis has been developed, which includes a scale-aware detection head, a feature fusion strategy, and the utilization of Online Hard Example Mining (OHEM) to improve detection performance while reducing computation overhead and addressing class imbalances in a real corpus dataset.
BNU-LCSAD[16] is a comprehensive dataset that can be employed for recognizing, detecting, and captioning students’ behaviors in a classroom. The dataset includes 128 videos from different disciplines across 11 classrooms. For different tasks, the baseline models used include the two-stream network for recognition, ACAM[17] and MOC[18] for action detection, BSN[19] and DBG[20] for temporal detection, and RecNet[21] and HACA[22] for video captioning.
Student Classroom Behavior Dataset[23] is a collection that includes 400 students from 90 classroom videos in a primary school. The dataset comprises single-person images, totaling 10,000 images of students engaging in behaviors such as raising their hands, walking back and forth, writing on the blackboard, and looking up and down, as well as an additional 1,000 images capturing students bending down, standing, and lying on the table. To recognize these classroom behaviors, a proposed method employs a 10-layer deep convolutional neural network (CNN-10) that extracts key information from human skeleton data, effectively excluding irrelevant information, and achieves higher recognition accuracy and generalization ability.
Student behavior dataset[24] is an intelligent classroom-based dataset that categorizes student behavior into seven classes. Challenging class surveillance videos were carefully selected and annotated to create this dataset. The proposed method within this context integrates relational features to analyze how actors interact with their surrounding context. It models human-to-human relationships using body parts and context, and then combines these relational features with appearance features to achieve accurate human-to-human interaction recognition.
Student action dataset[25] is a collection that comprises 3881 labeled images depicting a variety of student behaviors within a classroom setting, such as raising hands, paying attention, eating, being distracted, reading books, using phones, writing, feeling bored, and laughing. For the training and evaluation of these behaviors, the YOLOv5 object detection model is employed.
A large-scale student behavior dataset[26] contains five representative student behaviors highly correlated with student engagement, including raising hand, standing up, sleeping, yawning, and smiling, and tracks the change trends of these behaviors throughout the course. The proposed system, StuArt, is an innovative automated solution that enables instructors to closely monitor the learning progress of each student in the classroom. Additionally, StuArt includes user-friendly visualizations to facilitate instructors’ understanding of individual and overall learning status.
Classroom behavior dataset[27] consists of genuine images sourced from publicly available education videos that span primary, middle, and university levels. Comprising 4432 images and 151574 annotations, this dataset captures 7 common student behaviors—writing, reading, listening, raising hand, turning around, standing, and discussing—as well as 1 typical teacher behavior, which is guiding. To enhance teaching quality and provide feedback through real-time analysis of student behavior in the classroom, the authors propose BiTNet, a real-time object detection network. BiTNet is designed to address challenges that current methods face, such as occlusion and the detection of small objects in images.
Student Teacher Behavior Dataset (STBD-08)[28] is a comprehensive collection that contains 4432 images with 151574 labeled anchors, covering eight typical classroom behaviors. To address challenges such as occlusions, pose variations, and inconsistent target scales, the authors propose an advanced single-stage object detector known as ConvNeXt Block Prediction Head Network (CBPH-Net). This network is designed to effectively handle the complexities associated with detecting and analyzing teacher behaviors in classroom settings.
DBS Dataset[29] is a student classroom behavior dataset created for the context of smart education, comprising nine classes: listening, raising hands, standing up, reading, writing, looking around, lying on the desk, discussing, and other behaviors, with a total of 6890 annotated images. The proposed VWE-YOLOv8 algorithm enhances YOLOv8 with CSWin-Transformer for global feature extraction, LSKA for multi-scale recognition, SEAMHead for occlusion handling, and Slide Loss for sample imbalance, improving detection accuracy in classroom settings.
II-B Students classroom behavior detection
Mature object detection is used by more and more researchers in student behavior detection.
YAN Xing-ya[4] et al. proposed a classroom behavior recognition method that leverages deep learning. Specifically, they utilized the improved Yolov5 target detection algorithm to generate human detection proposals, and proposed the BetaPose lightweight pose recognition model, which is based on the Mobilenetv3 architecture, to enhance the accuracy of pose recognition in crowded scenarios.
ZHOU Ye[14] et al has proposed a method for detecting students’ behaviors in class by utilizing the Faster R-CNN detection framework. To overcome the challenges of detecting a wide range of object scales and the imbalance of data categories, the approach incorporates the feature pyramid and prime sample attention mechanisms.
Chen H [30] et al(using SCB-Dataset[1]) the classroom is advanced by an Improved YOLOv8 model. This method combines Res2Net with YOLOv8’s C2f module to form the C2f_Res2block for multi-scale feature extraction. It also integrates EMA for scale-sensitive detection and MHSA for enhanced feature representation, resulting in improved accuracy over the original YOLOv8.
Csb-yolo(using SCB-Dataset[1])[31] is a real-time algorithm for detecting classroom student behaviors. It uses BiFPN for feature fusion, an ERD Head for faster inference, and SCConv to maintain accuracy. The model is optimized with pruning and distillation for lightweight deployment on classroom devices.
III SCB-Dataset5
Classroom teaching has always played a fundamental role in education. Understanding students’ behavior is crucial for comprehending their learning process, personality, and psychological traits. In addition, it is an important factor in evaluating the quality of education. In our continuous practice and feedback, we have collected and annotated 20 types of student-teacher classroom behaviors (including 2 non-behavioral classes, blackboard and screen).
Our SCB-Dataset5 is divided into three major categories: student behavior, teacher behavior, and blackboard/screen. such as Table I
TABLE I: Classification of Behaviors in SCB-Dataset5
Student Behavior Individual Behavior hand-raising, answer, stage interaction, yawn, leaning on the desk, using the phone talk stand
Group Behavior read, write, bow head, turn head, discuss, clap, using the computer
Teacher Behavior guide, board writing, teacher
Blackboard -Screen blackboard, screen
In classroom settings, we believe that analyzing student behavior should be approached from two perspectives: group and individual.
Hand-raising is considered an individual behavior, where one or more students raise their hands to answer questions when the teacher asks. This behavior can reflect the teacher’s questioning style, teacher-student interaction, and the level of classroom activity.
We categorize behaviors such as reading and writing as group behaviors because observing individual students’ actions has limited significance. Instead, focusing on the general activities of students in the entire classroom is more important. The reading and writing behaviors of a single or a few students are not representative, while the majority of students being engaged in reading or writing can better reflect the current teaching state of the classroom.
Speaking and standing are behaviors that can belong to both teachers and students, so we classify them as shared behavior categories for teachers and students.
In addition to teacher and student behaviors, we have also included annotated data for blackboards and screens because, throughout the classroom, the content of the teacher’s writing on the blackboard and the content of the slides on the screen are very important for understanding the course material.
Below is the specific description table for the 20 types of behaviors, as shown in Table II.
TABLE II: SCB-Dataset5’s 20 Class Descriptions
Class Description
1 hand- raising Students actively raise their hands in class to indicate they want to speak or ask a question.
2 read Students read books, textbooks, or notes in class.
3 write Students take notes or complete written assignments in class.
4 bow head Students lower their heads to look at the desk or items in their hands, possibly being distracted or focused on personal activities.
5 turn head Students turn their heads, possibly to look at classmates/teachers or events happening in the classroom.
6 talk Students talk to others or answer questions in class.
7 guide Teachers provide guidance or explanations to students in class.
8 board writing Teachers/students write on the blackboard or draw.
9 stand Students or teachers stand in class.
10 answer Students respond to the teacher’s questions or instructions.
11 stage interaction Students interact with teachers or other students on the stage.
12 discuss Students discuss classroom content with each other.
13 clap Students applaud in class for a performance or achievement.
14 yawn Students yawn in class, possibly indicating tiredness or disinterest.
15 screen The electronic screen (such as a projector or computer screen) displays teaching content. Mainly used to capture the teacher’s courseware during class.
16 blackboard A traditional teaching aid, capturing key content from the teacher’s lecture. Mainly used to capture the teacher’s writing on the blackboard.
17 teacher The teacher responsible for teaching and guidance in the classroom, used to distinguish between student behavior and teacher behavior, such as whether it is the teacher writing on the blackboard or a student, or whether it is currently the teacher lecturing or a student answering a question.
18 leaning on the desk Students lean on their desks, possibly indicating tiredness or non-participation in classroom activities. Or it could be quiet rest before class.
19 using the phone Students use their phones in class, possibly being distracted or engaging in activities unrelated to the class. More common in universities.
20 using the computer Students use computers in class, more common in universities.
In reality, student and teacher’s behavior is often multifaceted and abundant. To capture this complexity, we directly collected real classroom recordings from websites like bjyhjy, 1s1k, and TikTok. Notably, we have also incorporated classroom data from China’s ethnic minorities. By using real-world videos, we ensured that our dataset is representative of real classroom situations, providing a more realistic and accurate reflection of student and teacher’s behavior.
Figure 3: Challenges in the SCB-Dataset5 include pixel differences, dense environments, the coexistence of multiple classes , and similar behaviors.
Classrooms are densely populated environments, and the SCB-Dataset5 also poses many challenges. For example, as shown in Fig 3 I, there is a significant pixel difference between the images of students in the front row and those in the back row. As shown in Fig 3 II, a classroom may have more than 100 students present at the same time. As shown in Fig 3 III, the same student/teacher may exhibit a variety of behaviors, such as standing and talking, reading while raising their hand, and raising their hand while lying on the desk, which is referred to as multiple classes. As shown in Fig 3 IV, there is a high degree of similarity between behaviors, such as the similarity between placing a hand on the forehead and raising a hand, or the similarity between writing and reading.
Figure 4: The diversity of the SCB-Dataset5 includes varying shooting angles, class differences, different learning stages, and different ethnic groups.
The SCB-Dataset5 exhibits a rich diversity, as shown in Fig. 4 I, encompassing a variety of perspectives within classroom settings, including frontal, lateral, and back views. The same behavior can significantly differ when viewed from various angles, which increases the complexity of behavior detection. As demonstrated in Fig. 4 II, the dataset also includes a range of classroom environments and course types; for instance, computer courses are typically conducted in well-equipped computer labs, while English and other cultural courses are held in standard classrooms, and art courses might take place in orderly arranged rehearsal rooms. As presented in Fig. 4 III, the dataset covers students’ growth stages from kindergarten through university, and as shown in Fig. 4 IV, it includes the diversity of different ethnic backgrounds. This comprehensive consideration across ages, cultures, and environments provides a more thorough and in-depth data foundation for research.
As shown in Table III, we have made the behavior data listed in the table publicly available and will continue to update and expand the data. Since multiple behaviors may coexist for students/teachers at the same time, but the current multi-class object detection performance is not ideal, we have grouped non-conflicting behaviors into the same category. Therefore, we have divided the SCB-Dataset5 into eight subsets: A, B, C, D, E, F, G, and H, as follows:
SCB5-A includes: hand-raising, reading, writing.
SCB5-B includes: screen, blackboard.
SCB5-C includes: discussing.
SCB5-D includes: guiding, answering, stage interaction, board writing.
SCB5-E includes: standing.
SCB5-F includes: teacher.
SCB5-G includes: bowing head, turning head.
SCB5-H includes: talking.
For each subset of the SCB-Dataset5, we randomly partition the dataset based on video frames to form training and validation sets with a ratio of approximately 4:1. Due to the varying number of frames in each video, the specific quantity ratio of the training and validation sets in Table 3 is roughly 4:1.
TABLE III: Class quantity statistics for SCB-Dataset5
Dataset Images Annotation
SCB5-A Total: 6864 Total: 47372
Train: 5193 Train: 34524 hand-raising: 10538
read: 17539
write: 6447
Val: 1671 Val: 12848 hand-raising: 2915
read: 6539
write: 3394
SCB5-B Total: 2441 Total: 5560
Train: 1855 Train: 4113 screen: 1566
blackboard: 2547
Val: 586 Val: 1447 screen: 514
blackboard: 933
SCB5-C Total: 864 Total: 5392
Train: 605 Train: 3607 discuss: 3607
Val: 259 Val: 1785 discuss: 1785
SCB5-D Total: 5940 Total: 6186
Train: 5155 Train: 5362 guide: 1286
answer: 2547
stage interaction: 516
board writing: 793
Val: 785 Val: 824 guide: 146
answer: 475
stage interaction: 62
board writing: 141
SCB5-E Total: 7602 Total: 12969
Train: 5661 Train: 9670 stand: 9670
Val: 1941 Val: 3299 stand: 3299
SCB5-F Total: 7428 Total: 7727
Train: 5752 Train: 6047 teacher: 6047
Val: 1676 Val: 1680 teacher: 1680
SCB5-G Total: 2410 Total: 16118
Train: 1905 Train: 12365 bow-head: 4422
turn-head: 7943
Val: 505 Val: 3753 bow-head: 540
turn-head: 3213
SCB5-H Total: 4539 Total: 5506
Train: 3513 Train: 4184 talk: 4184
Val: 1026 Val: 1322 talk: 1322
Figure 5: SCB-Dataset5 Class Count Statistics
IV Experiment
IV-A Experimental Environment and Dataset
We conducted our experiments on an NVIDIA GeForce RTX 2080 Ti GPU with 11 GB of video memory, running Ubuntu 20.04.2 as the operating system. The code was implemented in Python 3.8 and we used PyTorch version 1.8.1 with CUDA version 10.1 for model training.
The dataset used in our experiments is SCB-dataset5, which we split into training, validation sets with a ratio of 4:1.
IV-B Model Training
To train the model, set the epoch range from 30 to 50., batch size to 4, and image size to 640x640, and we use a pre-trained model for the training.
IV-C Experimental Results and Analysis
For our training, we employed YOLOv7 network structures. Our experimental results are shown in Table IV. , with precision represented by ”p”, recall represented by ”R”, mAP@0.5 represented mean Average Precision at Intersection over Union threshold of 0.5, mAP@0.95 represented mean Average Precision at Intersection over Union threshold of 0.95. The data units in the table are ”%”. In Table IV, due to the unsatisfactory training results from the ”bow head” and ”turn head”, we will not present the results.
TABLE IV: The training results of SCB-Dataset5 on YOLOv7.
Dataset class P R mAP@0.5 mAP@.95
SCB5-A all 71.1 70.9 74.0 56.8
hand-raising 79.4 76.9 79.2 59.4
read 65.5 68.2 70.5 52.9
write 68.4 67.8 72.2 58.1
SCB5-B all 94.5 97.3 98.7 91.9
screen 94.8 95.7 98.2 95.1
backboard 94.2 98.9 99.2 88.8
SCB5-C all/discuss 67.5 72.5 74.7 39.3
SCB5-D all 85.5 82.6 86.4 67.2
guide 88.0 81.7 87.0 49.5
answer 89.3 88.0 92.3 76.7
stage interaction 69.9 65.2 68.5 54.7
board writing 94.5 95.6 97.7 87.9
SCB5-E all/stand 95.8 91.7 96.6 80.5
SCB5-F all/teacher 96.2 94.4 97.7 82.7
SCB5-G all
bow-head - - - -
turn-head - - - -
SCB5-H all/talk 87.8 62.6 77.2 61.3
V Conclusion
In summary, this paper has made contributions to the field of student behavior detection in education. Through the development of the SCB-dataset and its evaluation using the YOLOv7 algorithm, we have addressed the gaps in student data sets within the field and provided fundamental data for future research. These contributions have the potential to enhance the accuracy and effectiveness of student behavior detection systems, ultimately benefiting both students and educators.
Despite expanding the scale of SCB-Dataset to make it more comprehensive, in real classroom settings, apart from behaviors with larger data volumes such as reading, writing, and hand-raising that perform well, the performance of other categories with smaller data volumes is not satisfactory.
We believe that our work will help to advance the application of artificial intelligence in education, enhance teaching effectiveness, and ultimately benefit students. Our SCB-dataset can be downloaded from the link provided in the abstract for further research and development.